A vacuum cleaner is plugged into a 120 V (rms) wall outlet and uses an average power of 3,150 W. If the power company charges $0.18 per kWh, how much does it cost to run the vacuum cleaner for 45 minutes?
Answer #1

Solution) Vrms = 120 V

P = 3150 W = 3.150 KW

Cost = $0.18 per KWh

Cost for 45 minutes = ?

1 h = 60 minutes

? h = 45 minutes

Time = (45/60) = (3/4) h

For 1 KWh = $ 0.18

Energy = Power×Time = 3.150×(3/4) = 2.3625 KWh

So cost for 2.3625 KWh = $0.18×2.3625 = $0.425

Therefore, Cost = $0.425